# RALL-E: Robust Codec Language Modeling with Chain-of-Thought Prompting for Text-to-Speech Synthesis

### v2

@@ Abstract
We present RALL-E, a robust language modeling method for text-to-speech (TTS) synthesis. While previous work based on large language models (LLMs) shows impressive performance on zero-shot TTS, such methods often suffer from poor robustness, such as unstable prosody (weird pitch and rhythm/duration) and a high word error rate (WER), due to the autoregressive prediction style of language models. The core idea behind RALL-E is chain-of-thought (CoT) prompting, which decomposes the task into simpler steps to enhance the robustness of LLM-based TTS. To accomplish this idea, RALL-E first predicts prosody features (pitch and duration) of the input text and uses them as intermediate conditions to predict speech tokens in a CoT style. Second, RALL-E utilizes the predicted duration prompt to guide the computing of self-attention weights in Transformer to enforce the model to focus on the corresponding phonemes and prosody features when predicting speech tokens. Results of comprehensive objective and subjective evaluations demonstrate that, compared to a powerful baseline method VALL-E, RALL-E significantly improves the WER of zero-shot TTS from $5.6\%$ (without reranking) and $1.7\%$ (with reranking) to $2.8\%$ and $1.0\%$, respectively. Furthermore, we demonstrate that RALL-E correctly synthesizes sentences that are hard for VALL-E and reduces the error rate from $68\%$ to $4\%$.
@@ 1 Introduction
Large language models (LLMs) have demonstrated great progress in natural language generation (Radford et al., 2019; Brown et al., 2020). With a sufficient model size LLMs emerge powerful in-context learning abilities that can handle unseen tasks with a text instruction (usually called prompt) in a zero-shot or few-shot manner (Wei et al., 2022a). Moreover, the simple yet effective next-token prediction task of LLMs makes it easy to apply LLMs on other fields, such as vision (Dehghani et al., 2023) and speech synthesis (Wang et al., 2023a), as long as the data can be converted to discrete speech tokens. This work focuses on the language modeling of text-to-speech (TTS) synthesis. Recent work (Wang et al., 2023a; Kharitonov et al., 2023) have shown that TTS can be modeled by a decoder-only language model by using a neural codec (Zeghidour et al., 2021; Défossez et al., 2022) to convert continuous waveforms into discrete tokens. These methods, typically leverage tens of thousands of hours of speech data, emerge in-context learning ability that can clone a speaker’s voice by providing a short audio prompt to the language model, and thus have impressive performance on zero-shot TTS.
Table 1: Performance of RALL-E and the baseline method VALL-E (Wang et al., 2023a) on $50$ particularly hard sentences obtained from Ren et al. (2019). The result of NaturalSpeech 2 is from Shen et al. (2023).
@@
However, due to the sequential generation property of language models, such codec language models suffer from poor robustness. Although the autoregressive (AR) prediction style of language models enables the model to generate speech with diverse prosody patterns, they can also cause bad cases with unnatural prosody. Moreover, since there is no strict alignment between text and speech, the models can omit or repeat words in the input text. This is quite different from TTS methods based on non-autoregressive (NAR) generative models (Shen et al., 2023; Le et al., 2024; Ju et al., 2024), which predicts all tokens at the same time, thus have high robustness but relatively low diversity. As suggested by previous work (Yang et al., 2023; Ju et al., 2024), LLM-based TTS have a higher word error rate (WER) than NAR TTS even if they have similar performance on other metrics. To alleviate this problem, a simple but effective method is to sample the same input text multiple times and select the best one (Kharitonov et al., 2023; Yang et al., 2023). However, such a reranking method further increases the inference time.
@@
In this paper, we present RALL-E (the abbreviation of robust VALL-E), a method to improve the robustness of LLM-based TTS. The core idea of RALL-E is inspired from the chain-of-thought (CoT) prompting (Wei et al., 2022b). In CoT prompting, the LLM is instructed to generate an intermediate result that is used as a condition for the prediction of the final result. The CoT prompting breaks a complex task into several simpler steps, so that can improve the robustness of LLMs, especially on hard tasks like arithmetic (Wei et al., 2022b). To adapt CoT prompting to LLM-based TTS, RALL-E predicts prosody tokens (pitch and duration) before predicting speech tokens to stabilize the prosody. Given an input sentence, RALL-E first predicts phoneme-level pitch and duration of the input, then predicts speech tokens conditioning on both the input phonemes and the predicted prosody tokens. Furthermore, RALL-E utilizes the predicted duration to mask irrelevant phonemes and prosody tokens when computing self-attention weights, so that the codec language model is enforced to concentrate on tokens around the phoneme and prosody token the speech token corresponds to. We use VALL-E (Wang et al., 2023a), a recent powerful LLM-based TTS method, as the baseline, and conduct experiments on a large dataset with $44$K hours speech data. Results of comprehensive objective and subjective evaluations demonstrate that RALL significantly improves the robustness of LLM-based TTS by reducing the WER on the LibriSpeech (Panayotov et al., 2015) test-clean set from $5.6\%$ (w/o reranking) and $1.7\%$ (with reranking) to $2.8\%$ and $1.0\%$, respectively. Furthermore, we evaluate RALL-E on $50$ particularly hard sentences. As demonstrated in Table 1, compared to VALL-E, RALL-E significantly reduces the error rate from $68\%$ to $4\%$ by eliminating almost all types of error, which demonstrates the superior robustness of RALL-E (see Section 4.4 for more details). The contributions of this work are summarized as follows:
@@ 2 Related work
Inspired by the success of LLMs (Radford et al., 2019; Brown et al., 2020), several recent works adopt language models to model TTS (Wang et al., 2023a; Yang et al., 2023; Kharitonov et al., 2023) and begin to use decoder-only architecture based on transformer (Vaswani et al., 2017). In such models, text and speech tokens are concatenated together and fed to a single transformer. The whole model is trained on a next-token prediction task like a language model. The LLM-based TTS systems are typically trained on tens of thousands of hours of speech data and have hundreds of millions of parameters, hence can leverage the emergent abilities of LLMs like in-context learning (Wei et al., 2022a) to enable zero-shot TTS (Wang et al., 2023a). Besides, recent works (Rubenstein et al., 2023; Wang et al., 2023b; Yang et al., 2023) have shown the decoder-only architecture can be used to learn multiple tasks, as the input and output are processed jointly by a language model, and the model can be signaled to generate results for different tasks by inputting pre-defined special tokens. RALL-E focuses on the robustness problem of LLM-based TTS.
@@
The robustness of AR TTS is a popular topic in the literature. For encoder-decoder AR TTS, several previous works enforce the attention weights to be monotonic (Zhang et al., 2018; He et al., 2019; Chen et al., 2020) that can effectively improve the robustness. In addition, Shen et al. (2020) proposed a non-attentive Tacotron, in which the attention module was replaced by a duration predictor to determine the alignment path before decoding. For decoder-only TTS, a key difference is that the attention weights are computed on text and context at the same time, hence the whole attention weights should not be monotonic. Song et al. (2024) proposed ELLA-V that interleaves the speech tokens with phonemes by inserting a phoneme token and a special EndOfPhone (EOP) token at the beginning and end of the speech tokens corresponding to the phoneme, respectively. While the inserted phoneme and the EOP token indicate the duration of each phoneme, such an implicit way entangles the prediction of speech tokens and duration together. RALL-E disentangles the predictions of duration and speech tokens by predicting the duration of all phonemes before the speech tokens, hence has higher controllability over the generation process. Du et al. (2024) proposed VALL-T that uses an unsupervised transducer loss (Graves, 2012) to implicitly model the duration of phonemes. Compared to RALL-E, although VALL-T doesn’t rely on external alignment tools during training, its training time is considerably decelerated since the transducer loss requires the model to perform a forward process for every phoneme. Besides, like ELLA-V, VALL-T also entangles the predictions of duration and speech tokens, thus has weaker controllability than RALL-E.
@@ 3.1 Preliminary: VALL-E
We inherit most symbols from the original paper of VALL-E (Wang et al., 2023a) for ease of reading. Readers are recommended to refer to the original paper for more details.
Generally, VALL-E is a decoder-only LLM-based TTS system that uses two Transformers (Vaswani et al., 2017) to predict speech tokens from the text. The speech tokens here are extracted from EnCodec (Défossez et al., 2022), a neural audio codec based on residual vector quantization (RVQ) (Zeghidour et al., 2021) that can convert continuous speech signal into discrete tokens. After predicting the discrete tokens, the waveforms can be reconstructed by feeding the tokens into the decoder of EnCodec. An RVQ typically contains $N$ quantization layers ($N=8$ in VALL-E), hence at each time step the encoded speech has $N$ tokens. Formally, given speech $\mathbf{y}$ and its transcription $\mathbf{x}$, the discrete speech token matrix $\mathbf{C}$ encoded by the codec has a shape of $T\times N$, where $T$ is the total time step. In addition to $\mathbf{x}$, to clone a speaker’s voice and utilize the in-context learning ability of LLMs, VALL-E receives a short prompt $\tilde{\mathbf{C}}^{T^{{}^{\prime}}\times N}$ as input before predicting $\mathbf{C}$. Hence, VALL-E models and maximizes the following distribution:
@@ 3.2 Prosody tokens as chain-of-thought prompts
One of the problems of LLM-based TTS is that it directly generates speech from phonemes with no restriction on the prosody, e.g. pitch, duration, etc, which usually results in speech with unstable prosody. A similar problem is also observed in Wei et al. (2022b) where the authors find LLMs cannot directly answer a complex question like arithmetic and propose CoT prompting to solve this problem. The idea of CoT prompting is breaking a complex task into several simpler tasks so that LLMs can utilize the intermediate results to reach the final answers. As shown in Wei et al. (2022b), by CoT prompting the correct rate of LLMs on complex tasks can be significantly improved. This motivates us to adapt CoT prompting to LLM-based TTS by generating intermediate prosody tokens before generating speech tokens to alleviate the robustness problem of LLM-based TTS. To incorporate pitch and duration in the AR Transformer of VALL-E, we first get the alignment between phonemes and speech tokens and extract the pitch value for each speech token. We then compute phoneme-level pitch value based on the duration and linearly quantize it to $M_{p}$ buckets. We define a maximal duration value $M_{d}$, and all duration values that exceed $M_{d}$ will be truncated to the maximum. RALL-E predicts the two prosody tokens before the speech tokens in a CoT style. Formally, assume $\mathbf{p},\mathbf{d}$ are the discrete pitch and duration sequences of the target speech tokens $\mathbf{C}$, $\tilde{\mathbf{p}},\tilde{\mathbf{d}}$ are the ones of the prompt $\tilde{\mathbf{C}}$, we model and maximize the following distribution:
@@ 3.3 Enhancing alignment with duration-guided masking
The general inference procedure follows VALL-E (Wang et al., 2023a) with two differences. First, before sampling the speech tokens $\mathbf{c}_{:,1}$ the prosody tokens $\mathbf{p}$ and $\mathbf{d}$ are sampled conditioning on the phoneme sequence $\mathbf{x}$ and acoustic prompt $\tilde{\mathbf{p}},\tilde{\mathbf{d}}$. Second, although normal language models depend on a special token <eos> as the stop condition, since we know the total duration $D=\sum_{t=1}^{L}d_{t}$, we propose a duration-guided inference method that forces the inference to stop at the $D$-th step. This method ensures no phoneme is omitted or repeated as it continues the inference if the <eos> token is predicted before the $D$-th step and stops at the right step as guided by the predicted duration.
@@ 4.1 Setup
We use the English subset of the multilingual LibriSpeech (MLS) (Pratap et al., 2020) corpus containing about $44$K hours of speech data from $5490$ speakers as the training data. The test-clean set of the LibriSpeech corpus Panayotov et al. (2015) is used as the test set. Following Wang et al. (2023a), we only select utterances with a length of $4$-second to $10$-second, which results in $1205$ utterances. For each utterance in the test set, we randomly select another one of the same speaker and use the first $3$s of it as the prompt. All speech data are in $16$ kHz. The transcriptions are converted into phonemes by a grapheme-to-phoneme tool (Sun et al., 2019). The frame-level pitch value is extracted by the WORLD vocoder (Morise et al., 2016). We get the alignments between the phoneme sequences and the speech tokens with our internal alignment tool. The maximal duration value $M_{d}$ is set to $32$. The phoneme-level pitch value is then computed based on the alignments. The number of buckets $M_{p}$ for pitch quantization is set to $256$.
@@
We use SoundStream (Zeghidour et al., 2021) as the speech codec to get speech tokens and decode waveforms from the tokens. The architecture follows the original work and the number of quantization layer $N$ of RVQ is set to $16$. The codec language model follows VALL-E (Wang et al., 2023a), where both the AR and NAR models are a $12$-layer Transformer (Vaswani et al., 2017) with $1024$-dimensional token embeddings, sinusoidal positional embeddings, $4096$-dimensional feed-forward layers, and a dropout rate of $0.1$. The window size $k$ is set to $1$ without explicit statements (see Appendix A for a detailed explanation of how we choose this value).
@@
The SoundStream codec is trained on $8$ NVIDIA V100 GPUs with a batch size of $200$ per GPU. We use Adam (Kingma and Ba, 2015) as the optimizer with a learning rate of $42e-$ 4. The model converges in about $440$K steps. The AR and NAR Transformers are trained separately on $16$ AMD MI200 GPUs with a batch size of $7000$ speech tokens per GPU. Adam is again used as the optimizer. The scheduled inverse square root learning rate is used with $30$K warm-up steps and a peak learning rate of $45e-$ 4. Both of the two Transformers converge in about $500$K steps.
We adopt nucleus sampling (Holtzman et al., 2019) as the sampling method of the AR Transformer. For the predicted probability distribution, nucleus sampling first selects a token set with the highest probability whose accumulative probability exceeds a hyperparameter $\rho$, and randomly samples from the set. Note that, $\rho$ can be different for the sampling of pitch, duration, and speech tokens, which results in $3$ hyperparameters: $\rho_{p}$, $\rho_{d}$, $\rho_{c}$ for pitch, duration, speech tokens, respectively. Without explicit statements we set $\rho_{p}=\rho_{d}=\rho_{c}=0.9$ in the following experiments. We select the token with the highest probability in the NAR Transformer without sampling.
@@
We use VALL-E (Wang et al., 2023a) as the baseline method. We implement VALL-E and train it on our training set. Besides, two previous works: VALL-T (Du et al., 2024) and ELLA-V (Song et al., 2024) are also compared with RALL-E. VALL-T is trained on the LibriTTS (Zen et al., 2019) corpus with $520$ hours of speech data, and we get $500$ synthesized samples selected from the test-clean set of LibriTTS from the authors. ELLA-V (Song et al., 2024) is trained on the training set of the LibriSpeech Panayotov et al. (2015) corpus with $960$ hours speech data. We request the authors to run ELLA-V on our test set and get $904$ samples. We don’t use the result of the original paper (Song et al., 2024) since the continual generation method adopted by ELLA-V can result in a significantly better WER than non-continual generation, as suggested by Wang et al. (2023a).
Table 2: Main results of RALL-E on the LibriSpeech test set with $1205$ utterances. Bold indicates the best score. The WER shown in the parenthesis is obtained from the HuBERT model used in the original VALL-E paper (Wang et al., 2023a). ${}^{\dagger}$ means the results of VALL-E trained on LibriLight. ${}^{\ddagger}$ means the results of VALL-E trained on the English subset of MLS. We get the results of VALL-T ($500$ samples) (Du et al., 2024) and ELLA-V ($904$ samples) (Song et al., 2024) from the authors. RALL-E ($904$) means the results are computed on the same test set with ELLA-V including $904$ samples.
@@
${}^{\dagger}$ ($5.9$) - - $0.58$
${}^{\ddagger}$ $5.6$ (6.3)6.3)6.3 ) $1.7$ $3.9$ $0.49$ $2.8$ $1.5$ $1.3$
ELLA-V $4.6$ ($6.5$) - $2.2$ $0.40$ $3.0$ $1.2$ $0.4$
VALL-T $3.9$ ($5.4$) - $\mathbf{4.0}$ $0.46$ $2.4$ $1.3$ $\mathbf{0.2}$
RALL-E (904) $2.7$ ($3.0$) $0.8$ $0.49$ $1.6$ $0.7$ $0.3$
RALL-E $\mathbf{2.8}$ ($\mathbf{3.1}$) $\mathbf{1.0}$ $\mathbf{4.0}$ $0.49$ $\mathbf{1.7}$ $\mathbf{0.8}$ $0.3$
@@
  • Word error rate (WER). We transcribe the synthesized samples by a large Conformer-based (Gulati et al., 2020) ASR model [^1] https://huggingface.co/nvidia/stt_en_conformer_transducer_xlarge that is trained on a large collection of speech corpora including LibriSpeech (Panayotov et al., 2015). The WER is then computed between the recognized and the ground truth (GT) transcriptions. Besides, we also report WERs computed from transcriptions recognized by a HuBERT [^2] https://huggingface.co/facebook/hubert-large-ls960-ft (Hsu et al., 2021) model trained on Libri-Light (Kahn et al., 2020) and finetuned on LibriSpeech (Panayotov et al., 2015). We regard the WERs obtained from the Conformer-based model as the primary scores since it results in better performance than the HuBERT model, though the HuBERT model is used in the original VALL-E paper (Wang et al., 2023a).
@@
  • UTMOS (Saeki et al., 2022), which is a powerful automatic speech quality assessment model that evaluates speech naturalness.
  • Speaker similarity (SIM) defined as the cosine similarity between the speaker embeddings of the prompt and the synthesized utterance. Following VALL-E (Wang et al., 2023a), we use the wavlm_large_finetune checkpoint of WavLM-TDNN [^3] https://github.com/microsoft/UniSpeech/tree/main/downstreams/speaker_verification, a speaker verification model based on WavLM (Chen et al., 2022), to extract speaker embeddings.

### v3

@@ Abstract
We present RALL-E, a robust language modeling method for text-to-speech (TTS) synthesis. While previous work based on large language models (LLMs) shows impressive performance on zero-shot TTS, such methods often suffer from poor robustness, such as unstable prosody (weird pitch and rhythm/duration) and a high word error rate (WER), due to the autoregressive prediction style of language models. The core idea behind RALL-E is chain-of-thought (CoT) prompting, which decomposes the task into simpler steps to enhance the robustness of LLM-based TTS. To accomplish this idea, RALL-E first predicts prosody features (pitch and duration) of the input text and uses them as intermediate conditions to predict speech tokens in a CoT style. Second, RALL-E utilizes the predicted duration prompt to guide the computing of self-attention weights in Transformer to enforce the model to focus on the corresponding phonemes and prosody features when predicting speech tokens. Results of comprehensive objective and subjective evaluations demonstrate that, compared to a powerful baseline method VALL-E, RALL-E significantly improves the WER of zero-shot TTS from $5.6\%$ (without reranking) and $1.7\%$ (with reranking) to $2.5\%$ and $1.0\%$, respectively. Furthermore, we demonstrate that RALL-E correctly synthesizes sentences that are hard for VALL-E and reduces the error rate from $68\%$ to $4\%$.
@@ 1 Introduction
Large language models (LLMs) have demonstrated great progress in natural language generation [20, 1]. With a sufficient model size LLMs emerge powerful in-context learning abilities that can handle unseen tasks with a text instruction (usually called prompt) in a zero-shot or few-shot manner [31]. Moreover, the simple yet effective next-token prediction task of LLMs makes it easy to apply LLMs on other fields, such as vision [5] and speech synthesis [29], as long as the data can be converted to discrete speech tokens. This work focuses on the language modeling of text-to-speech (TTS) synthesis. Recent work [29, 14] have shown that TTS can be modeled by a decoder-only language model by using a neural codec [34, 4] to convert continuous waveforms into discrete tokens. These methods, typically leverage tens of thousands of hours of speech data, emerge in-context learning ability that can clone a speaker’s voice by providing a short audio prompt to the language model, and thus have impressive performance on zero-shot TTS.
Table 1: Performance of RALL-E and the baseline method VALL-E [29] on $50$ particularly hard sentences obtained from Ren et al. [21]. The result of NaturalSpeech 2 is from Shen et al. [25].
@@
However, due to the sequential generation property of language models, such codec language models suffer from poor robustness. Although the autoregressive (AR) prediction style of language models enables the model to generate speech with diverse prosody patterns, they can also cause bad cases with unnatural prosody. Moreover, since there is no strict alignment between text and speech, the models can omit or repeat words in the input text. This is quite different from TTS methods based on non-autoregressive (NAR) generative models [25, 16, 12], which predicts all tokens at the same time, thus have high robustness but relatively low diversity. As suggested by previous work [33, 12], LLM-based TTS have a higher word error rate (WER) than NAR TTS even if they have similar performance on other metrics. To alleviate this problem, a simple but effective method is to sample the same input text multiple times and select the best one [14, 33]. However, such a reranking method further increases the inference time.
@@
In this paper, we present RALL-E (the abbreviation of robust VALL-E), a method to improve the robustness of LLM-based TTS. The core idea of RALL-E is inspired from the chain-of-thought (CoT) prompting [32]. In CoT prompting, the LLM is instructed to generate an intermediate result that is used as a condition for the prediction of the final result. The CoT prompting breaks a complex task into several simpler steps, so that can improve the robustness of LLMs, especially on hard tasks like arithmetic [32]. To adapt CoT prompting to LLM-based TTS, RALL-E predicts prosody tokens (pitch and duration) before predicting speech tokens to stabilize the prosody. Given an input sentence, RALL-E first predicts phoneme-level pitch and duration of the input, then predicts speech tokens conditioning on both the input phonemes and the predicted prosody tokens. Furthermore, RALL-E utilizes the predicted duration to mask irrelevant phonemes and prosody tokens when computing self-attention weights, so that the codec language model is enforced to concentrate on tokens around the phoneme and prosody token the speech token corresponds to. We use VALL-E [29], a recent powerful LLM-based TTS method, as the baseline, and conduct experiments on a large dataset with $44$K hours speech data. Results of comprehensive objective and subjective evaluations demonstrate that RALL significantly improves the robustness of LLM-based TTS by reducing the WER on the LibriSpeech [18] test-clean set from $5.6\%$ (w/o reranking) and $1.7\%$ (with reranking) to $2.5\%$ and $1.0\%$, respectively. Furthermore, we evaluate RALL-E on $50$ particularly hard sentences. As demonstrated in Table 1, compared to VALL-E, RALL-E significantly reduces the error rate from $68\%$ to $4\%$ by eliminating almost all types of error, which demonstrates the superior robustness of RALL-E (see Section 4.4 for more details). The contributions of this work are summarized as follows:
@@ 2 Related work
Inspired by the success of LLMs [20, 1], several recent works adopt language models to model TTS [29, 33, 14] and begin to use decoder-only architecture based on transformer [28]. In such models, text and speech tokens are concatenated together and fed to a single transformer. The whole model is trained on a next-token prediction task like a language model. The LLM-based TTS systems are typically trained on tens of thousands of hours of speech data and have hundreds of millions of parameters, hence can leverage the emergent abilities of LLMs like in-context learning [31] to enable zero-shot TTS [29]. Besides, recent works [22, 30, 33] have shown the decoder-only architecture can be used to learn multiple tasks, as the input and output are processed jointly by a language model, and the model can be signaled to generate results for different tasks by inputting pre-defined special tokens. RALL-E focuses on the robustness problem of LLM-based TTS.
@@
The robustness of AR TTS is a popular topic in the literature. For encoder-decoder AR TTS, several previous works enforce the attention weights to be monotonic [36, 9, 2] that can effectively improve the robustness. In addition, Shen et al. [24] proposed a non-attentive Tacotron, in which the attention module was replaced by a duration predictor to determine the alignment path before decoding. For decoder-only TTS, a key difference is that the attention weights are computed on text and context at the same time, hence the whole attention weights should not be monotonic. Song et al. [26] proposed ELLA-V that interleaves the speech tokens with phonemes by inserting a phoneme token and a special EndOfPhone (EOP) token at the beginning and end of the speech tokens corresponding to the phoneme, respectively. While the inserted phoneme and the EOP token indicate the duration of each phoneme, such an implicit way entangles the prediction of speech tokens and duration together. RALL-E disentangles the predictions of duration and speech tokens by predicting the duration of all phonemes before the speech tokens, hence has higher controllability over the generation process. Du et al. [6] proposed VALL-T that uses an unsupervised transducer loss [7] to implicitly model the duration of phonemes. Compared to RALL-E, although VALL-T doesn’t rely on external alignment tools during training, its training time is considerably decelerated since the transducer loss requires the model to perform a forward process for every phoneme. Besides, like ELLA-V, VALL-T also entangles the predictions of duration and speech tokens, thus has weaker controllability than RALL-E.
@@ 3.1 Preliminary: VALL-E
We inherit most symbols from the original paper of VALL-E [29] for ease of reading. Readers are recommended to refer to the original paper for more details.
Generally, VALL-E is a decoder-only LLM-based TTS system that uses two Transformers [28] to predict speech tokens from the text. The speech tokens here are extracted from EnCodec [4], a neural audio codec based on residual vector quantization (RVQ) [34] that can convert continuous speech signal into discrete tokens. After predicting the discrete tokens, the waveforms can be reconstructed by feeding the tokens into the decoder of EnCodec. An RVQ typically contains $N$ quantization layers ($N=8$ in VALL-E), hence at each time step the encoded speech has $N$ tokens. Formally, given speech $\mathbf{y}$ and its transcription $\mathbf{x}$, the discrete speech token matrix $\mathbf{C}$ encoded by the codec has a shape of $T\times N$, where $T$ is the total time step. In addition to $\mathbf{x}$, to clone a speaker’s voice and utilize the in-context learning ability of LLMs, VALL-E receives a short prompt $\tilde{\mathbf{C}}^{T^{{}^{\prime}}\times N}$ as input before predicting $\mathbf{C}$. Hence, VALL-E models and maximizes the following distribution:
@@ 3.2 Prosody tokens as chain-of-thought prompts
One of the problems of LLM-based TTS is that it directly generates speech from phonemes with no restriction on the prosody, e.g. pitch, duration, etc, which usually results in speech with unstable prosody. A similar problem is also observed in Wei et al. [32] where the authors find LLMs cannot directly answer a complex question like arithmetic and propose CoT prompting to solve this problem. The idea of CoT prompting is breaking a complex task into several simpler tasks so that LLMs can utilize the intermediate results to reach the final answers. As shown in Wei et al. [32], by CoT prompting the correct rate of LLMs on complex tasks can be significantly improved. This motivates us to adapt CoT prompting to LLM-based TTS by generating intermediate prosody tokens before generating speech tokens to alleviate the robustness problem of LLM-based TTS. To incorporate pitch and duration in the AR Transformer of VALL-E, we first get the alignment between phonemes and speech tokens and extract the pitch value for each speech token. We then compute phoneme-level pitch value based on the duration and linearly quantize it to $M_{p}$ buckets. We define a maximal duration value $M_{d}$, and all duration values that exceed $M_{d}$ will be truncated to the maximum. RALL-E predicts the two prosody tokens before the speech tokens in a CoT style. Formally, assume $\mathbf{p},\mathbf{d}$ are the discrete pitch and duration sequences of the target speech tokens $\mathbf{C}$, $\tilde{\mathbf{p}},\tilde{\mathbf{d}}$ are the ones of the prompt $\tilde{\mathbf{C}}$, we model and maximize the following distribution:
@@ 3.3 Enhancing alignment with duration-guided masking
The general inference procedure follows VALL-E [29] with two differences. First, before sampling the speech tokens $\mathbf{c}_{:,1}$ the prosody tokens $\mathbf{p}$ and $\mathbf{d}$ are sampled conditioning on the phoneme sequence $\mathbf{x}$ and acoustic prompt $\tilde{\mathbf{p}},\tilde{\mathbf{d}}$. Second, although normal language models depend on a special token <eos> as the stop condition, since we know the total duration $D=\sum_{t=1}^{L}d_{t}$, we propose a duration-guided inference method that forces the inference to stop at the $D$-th step. This method ensures no phoneme is omitted or repeated as it continues the inference if the <eos> token is predicted before the $D$-th step and stops at the right step as guided by the predicted duration.
@@ 4.1 Setup
We use the English subset of the multilingual LibriSpeech (MLS) [19] corpus containing about $44$K hours of speech data from $5490$ speakers as the training data. The test-clean set of the LibriSpeech corpus [18] is used as the test set. Following Wang et al. [29], we only select utterances with a length of $4$-second to $10$-second, which results in $1205$ utterances. For each utterance in the test set, we randomly select another one of the same speaker and use the first $3$s of it as the prompt. All speech data are in $16$ kHz. The transcriptions are converted into phonemes by a grapheme-to-phoneme tool [27]. The frame-level pitch value is extracted by the WORLD vocoder [17]. We get the alignments between the phoneme sequences and the speech tokens with our internal alignment tool. The maximal duration value $M_{d}$ is set to $32$. The phoneme-level pitch value is then computed based on the alignments. The number of buckets $M_{p}$ for pitch quantization is set to $256$.
@@
We use SoundStream [34] as the speech codec to get speech tokens and decode waveforms from the tokens. The architecture follows the original work and the number of quantization layer $N$ of RVQ is set to $16$. The codec language model follows VALL-E [29], where both the AR and NAR models are a $12$-layer Transformer [28] with $1024$-dimensional token embeddings, sinusoidal positional embeddings, $4096$-dimensional feed-forward layers, and a dropout rate of $0.1$. The window size $k$ is set to $1$ without explicit statements (see Appendix A for a detailed explanation of how we choose this value).
@@
[…]
We adopt nucleus sampling [10] as the sampling method of the AR Transformer. For the predicted probability distribution, nucleus sampling first selects a token set with the highest probability whose accumulative probability exceeds a hyperparameter $\rho$, and randomly samples from the set. Note that, $\rho$ can be different for the sampling of pitch, duration, and speech tokens, which results in $3$ hyperparameters: $\rho_{p}$, $\rho_{d}$, $\rho_{c}$ for pitch, duration, speech tokens, respectively. Without explicit statements we set $\rho_{p}=\rho_{d}=\rho_{c}=0.9$ in the following experiments. We select the token with the highest probability in the NAR Transformer without sampling.
@@
We use VALL-E [29] as the baseline method. We implement VALL-E and train it on our training set. Besides, two previous works: VALL-T [6] and ELLA-V [26] are also compared with RALL-E. VALL-T is trained on the LibriTTS [35] corpus with $520$ hours of speech data, and we get $500$ synthesized samples selected from the test-clean set of LibriTTS from the authors. ELLA-V [26] is trained on the training set of the LibriSpeech [18] corpus with $960$ hours speech data. We request the authors to run ELLA-V on our test set and get $912$ samples. We don’t use the result of the original paper [26] since the continual generation method adopted by ELLA-V can result in a significantly better WER than non-continual generation, as suggested by Wang et al. [29].
Table 2: Main results of RALL-E on the LibriSpeech test set with $1205$ utterances. Bold indicates the best score. The WER shown in the parenthesis is obtained from the HuBERT model used in the original VALL-E paper [29]. means the results of VALL-E trained on LibriLight. means the results of VALL-E trained on the English subset of MLS. We get the results of VALL-T ($500$ samples) [6] and ELLA-V ($912$ samples) [26] from the authors. RALL-E ($912$) means the results are computed on the same test set with ELLA-V including $912$ samples.
@@
VALL-E - ($5.9$) - - $0.58$
VALL-E $5.6$ (6.3)6.3)6.3 ) $1.7$ $3.9$ $0.49$ $2.8$ ($3.6$) $1.5$ ($1.4$) $1.3$ ($1.3$)
ELLA-V ($912$) $2.8$ ($4.1$) $0.8$ $3.7$ $0.42$ $2.2$ ($3.4$) $0.4$ ($0.4$) $0.2$ ($0.3$)
VALL-T ($500$) $3.9$ ($5.4$) - $\mathbf{4.0}$ $0.46$ $2.4$ ($3.6$) $1.3$ ($1.6$) $\mathbf{0.2}$ ($0.2$)
RALL-E ($912$) $2.3$ ($2.6$) $0.8$ $4.0$ $0.49$ $1.4$ ($2.0$) $0.6$ ($0.3$) $0.3$ ($0.3$)
RALL-E $\mathbf{2.5}$ ($\mathbf{2.8}$) $\mathbf{1.0}$ $\mathbf{4.0}$ $0.49$ $\mathbf{1.7}$ ($2.2$) $0.6$ ($0.3$) $0.3$ ($0.3$)
@@
  • Word error rate (WER). We transcribe the synthesized samples by a large Conformer-based [8] ASR model [^1] https://huggingface.co/nvidia/stt_en_conformer_transducer_xlarge that is trained on a large collection of speech corpora including LibriSpeech [18]. The WER is then computed between the recognized and the ground truth (GT) transcriptions. Besides, we also report WERs computed from transcriptions recognized by a HuBERT [^2] https://huggingface.co/facebook/hubert-large-ls960-ft [11] model trained on Libri-Light [13] and finetuned on LibriSpeech [18]. We regard the WERs obtained from the Conformer-based model as the primary scores since it results in better performance than the HuBERT model, though the HuBERT model is used in the original VALL-E paper [29].
@@
  • UTMOS [23], which is a powerful automatic speech quality assessment model that evaluates speech naturalness.
  • Speaker similarity (SIM) defined as the cosine similarity between the speaker embeddings of the prompt and the synthesized utterance. Following VALL-E [29], we use the wavlm_large_finetune checkpoint of WavLM-TDNN [^3] https://github.com/microsoft/UniSpeech/tree/main/downstreams/speaker_verification, a speaker verification model based on WavLM [3], to extract speaker embeddings.